OpenStudy (anonymous):

metal and fan. 8th grade pre- algebra

OpenStudy (anonymous):

Paula and Mario are part of a chili cook-off. Each person needs to serve 14 children and 28 adults. Paula has already spent $31.56 on all the ingredients but the beef. Mario has already spent $37.60 on all the ingredients but the beef. Paula is now buying beef for $2.25 per pound. Mario is now buying beef for $1.25 per pound. How many pounds of beef, p, will Paula and Mario need to buy until they have spent the same total amount of money? Which equation below correctly represents the situation above?

OpenStudy (anonymous):

a. 31.56 + 37.6 + 42 = 2.25p + 1.25p b. 31.56 + 2.25p = 37.6 + 1.25p c.31.56 + 37.6 + 2.25 + 1.25 + 42 = p d. 2.25p = 1.25p

OpenStudy (anonymous):

31.56 + 2.25x = ? 37.60 + 1.25x = ? we are tying to get the same amount, correct?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

alright, then, the basic facts are what i stated above. do any of the answer choices make sense to you?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

you got it!